OpenStudy (anonymous):

the function f(x)= x^4+cos^2x is a) an even function b) odd function c) neither even nor odd d) a constsnt term

OpenStudy (jack12):

Its E, a cheesecake

terenzreignz (terenzreignz):

What do your instincts tell you? If they're mum on the matter, try evaluating f(-x) like last time... :D

OpenStudy (binary3i):

do the same thing put -x at the place of x. and check it f(-x)=f(x) its an even And if f(-x)=-f(x) its an odd

OpenStudy (jack12):

First factor out the frosting, then eat the jelly

ganeshie8 (ganeshie8):

also, cos(-x) = cos(x) may help u here

OpenStudy (anonymous):

ok its even !

OpenStudy (jack12):

no its odd

ganeshie8 (ganeshie8):

even = \(\checkmark\)

OpenStudy (anonymous):

it is even cause 4 times minus is +

OpenStudy (anonymous):

also with the cos thing

ganeshie8 (ganeshie8):

yep